### v1

# Applications of higher order Markov models and Pressure Index to strategize controlled run chases in Twenty20 cricket

Abstract
In limited overs cricket, the team batting first posts a target score for the team batting second to achieve in order to win the match. The team batting second is constrained by decreasing resources in terms of number of balls left and number of wickets in hand in the process of reaching the target as the second innings progresses. The Pressure Index, a measure created by researchers in the past, serves as a tool for quantifying the level of pressure that a team batting second encounters in limited overs cricket. Through a ball-by-ball analysis of the second innings, it reveals how effectively the team batting second in a limited-over game proceeds towards their target. This research employs higher order Markov chains to examine the strategies employed by successful teams during run chases in Twenty20 matches. By studying the trends in successful run chases spanning over $16$ years and utilizing a significant dataset of $3378$ Twenty20 matches, specific strategies are identified. Consequently, an efficient approach to successful run chases in Twenty20 cricket is formulated, effectively limiting the Pressure Index to $[0.5,3.5]$ or even further down under $0.5$ as early as possible. The innovative methodology adopted in this research offers valuable insights for cricket teams looking to enhance their performance in run chases.
Keywords: Cricket analytics, Data mining in sports, Markov chain, Markov model, Pressure Index, Transition matrix
1 Introduction
Cricket is a team sport played between two teams consisting of 11 players each, following a set of rules as per the format of the game being Tests, One Day Internationals or Twenty20 Internationals (First Class, List A and Twenty20 respectively on a larger basis). In addition, shorter formats of the game like T10, The Hundred, etc., are being introduced for rapid spread and popularity of the game worldwide. One of the major reasons for the intense popularity of cricket nowadays is the uncertainty during a game and it has been observed that the shorter the format, the higher the uncertainty. The above claim is in par with the fact that in the $5$-year period from July $2014$ to June $2019$, $29.89\ \%$ Test matches have been won by teams outside the top $3$ ICC men’s Test rankings against teams ranked within the top $3$ ICC men’s Test rankings, which increases to $33.67\ \%$ if we consider Twenty20 cricket during the same period. The rise is even more in the following $5$-year period from July $2019$ to June $2024$ where it increases from $19.48\ \%$ in Tests to $29.30\ \%$ in Twenty20 cricket.  [^3] Source of data: https://stats.espncricinfo.com/ci/engine/stats/index.html
In Twenty20 cricket, a team bats for a maximum of 20 overs with 10 wickets in hand and tries to score as many runs as possible within the provided resources. The team which fields first then comes in to bat with identical resources, being set a target to score more runs than the team which batted first. Unless in rain-interrupted matches or other scenarios resulting in reduction of overs, the team which scores more runs wins the match. The match is called a tie in case both teams end up scoring equal number of runs. Super overs (earlier, bowl outs) are often played to decide the outcome of tied matches.
As the game progresses, efficient use of numbers, statistics and technology plays a key role to analyze and strategize the game from a player and team’s point of view. The present study is an endeavor to introduce a strategy whereby teams could constantly set a target, within themselves, for their batters at certain intervals throughout the innings while chasing a score in a regular Twenty20 match with the help of the Pressure Index and higher order Markov chains. The Pressure Index is a ball-by-ball measure which quantifies the pressure experienced by the team batting second. It is a function of the required run rate and wickets lost at any point in the second innings, for our purposes, in a Twenty20 match. Thus, during a run chase, the value of the Pressure Index might change after every ball. The value of the Pressure Index starts at unity at the start of a run chase and keeps on changing as the innings progresses. It eventually reaches zero if the target is achieved but rises steeply if the team batting second fails to achieve the target. Details on working formulae, functionality, etc., regarding the Pressure Index are discussed later. The curve obtained by plotting the Pressure Index values against each ball during an innings is said to be the Pressure Curve. The Pressure Curve gives a clear indication of the status of the batting team at any given instance during a run chase. In this paper, Pressure Index values have been obtained corresponding to numerous Twenty20 matches spanned over a prolonged period. Then the concepts of Markov chains have been used in a way to learn the pattern of changes in the Pressure Index throughout a run chase in Twenty20 cricket and an attempt has been made to suggest a suitable strategy to keep the Pressure Index under control so that teams do not often lose grip on run chases.
2 Review of literature
Though the paper is heavily dependent on the Pressure Index for formulating a strategic approach towards successful run chases in the Twenty20 format of cricket, the work of [Bhattacharjee and Lemmer, 2016] and their subsequent extended works, to name a few, [Talukdar and Bhattacharjee, 2018] and [Bhattacharjee et al., 2018] are not the only contributions in this domain of tactic building for run chases.
Some earlier works that were concerned with the progress of scoring in the second innings of limited overs matches are [Bailey and Clarke, 2006] and [De Silva et al., 2001] While [Bailey and Clarke, 2006] attempted to predict the final score of the second innings of a One-day match while the game was in progress, [De Silva et al., 2001] attempted to estimate the magnitude of victory in One-day matches. In both cases the authors used the Duckworth–Lewis resource table at the end of each over of the second innings of the match.
[Choudhury and Bhargava, 2007] used Multilayer Perceptron Neural Networks to predict the outcome of a cricket tournament involving more than two teams. [Sankaranarayanan, 2014] developed a prediction system that takes in historical ODI cricket match data as well as the instantaneous state of a match and predicts the score at key points in the future, culminating in a prediction of victory or loss. Using factors of the current situation of a cricket match, [Lokhande and Chawan, 2018] attempted to predict the outcome through a two-stage process. In the first stage, the match situation and other cricketing attributes are used to predict the first innings score and then in the second stage the possibility of the team batting second achieving the target is estimated. [Asif and McHale, 2016] presented a model for forecasting the outcomes of One-Day International cricket matches when the game is in progress using dynamic logistic regression. [O’Riley and Ovens, 2006] compared four different score prediction tools by extending them to a ball-by-ball prediction model for predicting the final score, based on the present state of the first innings of a match. A series of papers are available in the literature that used different machine learning tools for predicting either the target of the match or the winner of the match when the match is in progress, viz. [Pathak and Wadhwa, 2016], [Passi and Pandey, 2018], [Vistro et al., 2019], [Modekurti, 2020]. Another set of researchers worked on the development of the Win and Score Predictor (WASP). The purpose of WASP is to predict the score at the end of the first innings of a limited overs match given the current situation and to predict the probability of a successful run chase of the team batting second given the target. Some literature concerning this includes [Hogan et al., 2017], [Singh et al., 2015], and [Nimmagadda et al., 2018].
3 Objectives
The paper achieves the following objectives:
  1. 1.
    To develop a Markov model to quantify whether teams have exhibited any set pattern in successful run chases over a prolonged period of time.
  2. 2.
    To study the patterns of such run chases and accordingly prepare a strategy in order to keep the team batting second in control throughout a run chase for future Twenty20 matches.
4 Concepts of Markov Chain
4.1 Markov chain
Definition 4.1.1  A random process $\{X_{n}:n\geq 0\}$ is said to be a Markov chain with state space $S$ if it satisfies the Markov condition, i.e., for any states $x_{0}$, $x_{1}$, $x_{2}$, …, $x_{n-1}$, $i$, $j\in S$,
$\mathbb{P}(X_{n+1}=j\mid X_{n}=i,X_{n-1}=x_{n-1},\dots,X_{1}=x_{1},X_{0}=x_{0})=\mathbb{P}(X_{n+1}=j\mid X_{n}=i)$ (1)
It can be easily verified that the Markov condition is equivalent to the condition: for any states $x_{1}$, $x_{2}$, …, $x_{k-1}$, $i$, $j\in S$,
$\mathbb{P}(X_{n+1}=j\mid X_{n_{k}}=i,X_{n_{k-1}}=x_{k-1},\dots,X_{n_{1}}=x_{1})=\mathbb{P}(X_{n+1}=j\mid X_{n_{k}}=i)$ (2)
for any $n_{1}<n_{2}<\ldots<n_{k}\leq n$.
The evolution of the chain clearly depends on the probabilities $\mathbb{P}(X_{n+1}=j\mid X_{n_{k}}=i)$. In general, this depends on $i$, $j$ and $n$. We restrict our attention to the case where this probability is independent of $n$.
Definition 4.1.2  A transition matrix $\mathbf{P}=((p_{ij}))$ is the matrix whose $(i,j)$th entry is given by
$((p_{ij}))=\mathbb{P}(X_{1}=j\mid X_{0}=i).$ (3)
It is clear that, being conditional probabilities, the $p_{ij}$ ’s must satisfy
$p_{ij}\geq 0\quad\text{and}\quad\sum_{j\in S}p_{ij}=1\quad\text{for all}\quad i\in S.$ (4)
4.2 Markov chains of higher order
The setup mentioned in Section 4.1, where the probability of an event at any time point depends only on that of the immediate past, can be thought of as a first order Markov chain. We may encounter situations where the probability of an event at a certain time point (say, the current state) might depend not just on the immediate past but on several time points prior to the current state. These are considered to be higher order Markov chains.
Definition 4.2.1  A random process $\{X_{n}:n\geq 0\}$ is said to be a Markov chain of order $k$ if for any states $x_{0}$, $x_{1}$, $x_{2}$, …, $x_{n-1}$, $i$, $j\in S$,
$\mathbb{P}(X_{n+1}=j\mid X_{n}=i,\dots,X_{0}=x_{0})=\mathbb{P}(X_{n+1}=j\mid X_{n}=i,\dots,X_{n-k+1}=x_{n-k+1}).$ (5)
For our purpose, consider $\{X_{n}:n\geq 1\}$ to be the sequence of Pressure Index values (detailed in Section 5) for a team batting second at the end of each over during a run chase in a Twenty20 match, i.e., $X_{n}$ is the Pressure Index value at the end of the $(n+1)$th over during a run chase in a Twenty20 match. We will see why such a sequence is a suitable choice for Markov modelling.
5 Introduction to Pressure Index
5.1 Pressure Index
As mentioned earlier, the Pressure Index (PI) is a ball-by-ball measure which quantifies the pressure experienced by the team batting second. In Twenty20 cricket, the team batting second has to achieve a certain target before its resources are exhausted. Thus, the pressure of the team batting second reduces if runs are scored at a fast rate and if a considerable number of wickets are kept in hand until the target is reached. Based on this idea, the formula of the Pressure Index [Bhattacharjee and Lemmer, 2016] for teams batting second in Twenty20 matches is given by:
$PI=\frac{CRRR}{IRRR}\times\frac{1}{2}\left[e^{\frac{RU}{100}}+e^{\frac{\sum w_{i}}{11}}\right]$ (6)
where $IRRR$ is the initial required run rate, i.e., if $T$ is the target to be scored by the team batting second in $B$ balls, then
$IRRR=\frac{T\times 6}{B}.$ (7)
$CRRR$ denotes the current required run rate at any point during the second innings. If $R^{\prime}$ runs have been scored in $B^{\prime}$ balls at some stage, then
$CRRR=\frac{(T-R^{\prime})\times 6}{B-B^{\prime}}.$ (8)
The ratio $\frac{CRRR}{IRRR}$ thus measures the progress of the team batting second during a run chase. A lower value of the ratio denotes that the team is supposedly ahead in the chase.
The term $\sum w_{i}$ denotes the sum of the weights of the wickets that have fallen until any stage of the innings. As wickets fall, the team’s wicket strength deteriorates. For example, if a top order batter gets out, the wicket strength of the team drops more than when a lower order batter gets out. In order to account for the varying abilities of batters affecting their wicket strengths, wicket weights ($w_{i}$) as per [Lemmer, 2015] are used instead of assigning equal weights to every wicket. The wicket weights for different batting positions, as assigned by [Lemmer, 2015], are given below.
Position ($i$) 1 2 3 4 5 6 7 8 9 10 11
Wicket weight ($w_{i}$) 1.30 1.35 1.40 1.45 1.38 1.18 0.98 0.79 0.59 0.39 0.19
Finally, $RU$ denotes the percentage of resources used by the team at any stage during the innings as per the Duckworth–Lewis table.
Several other formulae to calculate the Pressure Index are available but we will avoid using those in this study due to various limitations. One such formula, as suggested by [Shah and Shah, 2014], is as follows:
$PI=CI\times 100+\left[\frac{wicket\,weight}{180}\times T\times\frac{Br}{B}\times\frac{Rr}{T}\right]$ (9)
According to the above formula, the PI value will always be 100 at the start. After calculating PI values for a considerable number of matches using this formula, it was found that PI decreased when a wicket fell and when the number of runs scored in an over was less than the required run rate. This is completely unrealistic considering the very purpose of the formulation of the Pressure Index and hence we will not take into account this formula.
Two other formulae which were also suggested, but had significant limitations, are given below.
$PI_{1}=\frac{CRRR}{IRRR}\times e^{\frac{\sum w_{i}}{11}}$ (10)
$PI_{2}=\frac{CRRR}{IRRR}\times e^{\frac{RU}{100}}$ (11)
The factor $CI=\frac{CRRR}{IRRR}$ is expected to change after every ball. While $e^{\frac{\sum w_{i}}{11}}$ changes only if a wicket falls, $e^{\frac{RU}{100}}$ changes after every ball as $RU$ takes into consideration not only the number of wickets fallen but also the balls faced by a team. At the start of the second innings, the Pressure Index of a team is equal to 1 according to both $PI_{1}$ and $PI_{2}$ and it will fluctuate as the match progresses. If the team batting second is able to achieve the target, both $PI_{1}$ and $PI_{2}$ become 0. However, it seems that $PI_{2}$ is more sensitive compared to $PI_{1}$ because it depends on $RU$, which is expected to change after every ball throughout the innings.
Considering all these factors, the formula given in 6 seems to be the most commendable, and we will consider the same for further study in this paper.
5.2 Working of the Pressure Index
It is evident from 6 that the Pressure Index value of a team batting second remains at unity at the start of a run chase. With every delivery bowled, the value might change and it gradually reaches zero once the team has achieved its target. In case the team fails to do so, the PI value rises steeply at the later stages of the match. To better understand how the indices work, the example of two matches have been taken.
Consider the third Twenty20 match played between South Africa and India at Newlands Cricket Ground, Cape Town on February 24, 2018. South Africa started slowly but cautiously, while chasing 173, being $45/1$ after 9 overs. Although the required run rate was creeping up, 9 wickets were still in hand and they were still in the game. But just when they started to accelerate, wickets began falling, which resulted in a halt in their scoring rate leading to a rapid increase in the Pressure Index. By the 18th over, SA completely lost their grip on the run chase and finally fell short by 7 runs even after scoring 45 runs off their last 3 overs. On the other hand, Australia succesfully chased down a record 244 runs just a week earlier at Auckland with the strategy of never letting the required runrate rise too high.
Figure 1: Pressure Curves for T20I 649 (NZ v AUS 2018) and T20I 655 (SA v IND 2018)
Note: The dots in the Pressure curves denote fall of wickets
The urgency to score runs during a run chase increases as the game progresses, unless the target is too small to create any trouble for the team batting second. Also, scoring rates might halt when a wicket falls. The number of runs scored and wickets lost until any stage of the innings influence the scoring rate of the team in the next over. Furthermore, the target, the runs scored, and wickets lost until any stage of the innings determine the Pressure Index of the team at that stage. Thus, the Pressure Index at the end of an over during any stage of the innings depends on the Pressure Index of the team in the previous overs.
Thus, we can consider a sequence $\{X_{n}:n\geq 1\}$ where $X_{n}$ is the Pressure Index value at the end of the $(n+1)$th over during a run chase in a Twenty20 match. This leads to a setup where the concepts of Markov chains on the sequence $\{X_{n}:n\geq 1\}$ can be purposefully implemented.
6 Methodology
6.1 Estimating the order of Markov chain
We have defined our sequence to be the chronological $PI$ values at the end of each over of the second innings of a Twenty20 match but the number of preceding $PI$ values on which the $PI$ at a certain stage of the sequence depends is still unknown. In an ideal scenario, the runs scored and wickets lost in a certain over should technically depend on that in each of the previous overs in the innings. However, in order to avoid unnecessarily complex calculations and to achieve an optimum level of accuracy, one should try to find out the desired order of the Markov chain, i.e. the desired number of preceding $PI$ values on which the next $PI$ value in a sequence depends, that fits our data and leads to a high enough precision for prediction purposes.
All men’s Twenty20 matches in which the team batting second have either reached the target after batting for more than 18 overs or failed to reach the target and have lost the match by a margin not more than 10 runs have been considered as the training sample for the study. Since the goal of our study is to prepare a strategy in support of teams batting second in Twenty20 matches, a certain section of the selected sample of matches will be helpful in replicating the patterns of successful run chases. The choice of such selection has also been made based on the idea of using a significantly large training sample in order to enhance the construction of the model.
Starting from the first ever officially recorded men’s Twenty20 match on 13th June 2003 till 2nd June 2024, a total of 4775 matches have been played where the outcomes have been either of the ones mentioned earlier. After removing matches containing missing data and those in which complete ball-by-ball commentary are unavailable, we finally take into account a filtered set of 3778 matches. All relevant data regarding these matches are collected from https://www.espncricinfo.com to create a set of 3778 sequences of the form $\{X_{n}:n\geq 1\}$. The correct order of the Markov chain is then estimated using these sequences to proceed with further study.
We know that for a Markov chain of order $k$, the current state depends on the previous $k$ states and the transition probabilities represent the probability of transitioning from one sequence of $k$ states to the immediate next state. With the help of our available data, likelihood functions are calculated corresponding to the cases when $k=1,2$ and $3$. It has been observed that Markov chains of order 4 or more would force us to deal with unnecessarily complex calculations without providing a significant rise in accuracy in predictive purposes.
Higher likelihood values indicate a model that fits the data better. However, the likelihood alone is not sufficient for model selection because it is a non decreasing function of the number of parameters. This may lead to overfitting, where a more complex model fits the training data very well but performs poorly on new or unseen data. To determine the order which fits best to the data, AIC (Akaike Information Criterion) and BIC (Bayesian Information Criterion) have been considered which balance goodness of fit along with model complexity.
AIC and BIC are formulated as follows:
$AIC=2m-2\ln(L)$ (12)
$BIC=m\ln(n)-2\ln(L)$ (13)
where
$m=$ number of parameters in the model,
$L=$ likelihood of the model,
$n=$ number of observations ($=3778$).
Note that the number of parameters is the number of independent transition probabilities that need to be estimated. This depends on the order of the Markov chain and the number of unique states in the data. It can be worked out that $m=s^{k}(s-1)$, where $s$ denotes the number of states and $k$ is the order of the Markov chain.
Models with the lowest $AIC$ and $BIC$ values are identified. If both $AIC$ and $BIC$ correspond to the same order, that is a strong indication of being the best order for the Markov chain. If $AIC$ and $BIC$ suggest different orders, decisions are made based on the specific context and purpose of the model. Since BIC is more stringent, it is often preferred in the case of a large dataset to avoid overfitting. Fortunately, in this case, the lowest $AIC$ and $BIC$ values point to the same order for every sequence, so we do not encounter any ambiguity.
After inserting the data and computing likelihood, AIC and BIC values, it is observed that successful calculations have been carried out for 2982 out of the 3778 sequences. More than $95\%$ of those 2982 sequences have yielded 3rd order Markov chains  [^4] For Markov chain modeling, we primarily used the ’markovchain’ package in R. This package provides an S4 class structure for representing Markov chain objects and includes functions (e.g., for fitting, simulating, and analyzing chains) that underpin our custom transition matrix construction and subsequent analysis of Pressure Index transitions. to be the best fitted estimate.
Final counts according to AIC:
Estimated order Number of sequences
1 144
2 2
3 2836
Final counts according to BIC:
Estimated order Number of sequences
1 144
2 2
3 2836
$95.10\%$ of the sequences result in a best fit of order 3 and as a result, we decide to proceed considering a Markov chain of order 3 for the overall data. A detailed reasoning for the same is as follows:
  1. i.
    Higher order dependencies: A higher-order Markov model captures more complex dependencies by considering a longer history of past states. This means an order-3 model takes into account the previous three states to predict the current state, whereas a lower-order model considers fewer past states.
  2. ii.
    Uniformity in model structure: Using the same order for all sequences ensures that the model structure is consistent across all sequences. This uniformity is crucial for combining data from different sequences while constructing a single transition matrix.
  3. iii.
    Avoiding underfitting: Using a lower order might underfit the data, missing out on important dependencies captured by higher order models.
Using the highest order ensures that the full extent of dependencies present in the sequences is captured, providing a more accurate and informative transition matrix. While this might be more complex, it avoids the risk of missing critical patterns that could have been obscured by using a lower order model.
6.2 Applying the Markov model on Pressure Index
From each sequence of the entire data, every possible state, say $x_{t}$, is considered in the columns as the current state and every set of preceding states in the form of a tuple $(x_{t-3},x_{t-2},x_{t-1})$ is considered in the rows as the previous states. Note that $\{X_{n}:n\geq 1\}$ is the sequence of Pressure Index values at the end of each over during the second innings of a match, and so, $x_{t}\in S$, where $S$ is the state space that contains all possible values the Pressure Index of a team can take. The transition probabilities $p_{ij}$ are then obtained from the data and the resulting matrix satisfies the properties mentioned in Definition 4.1.2. The obtained transition matrix  [^5] The transition matrix obtained using PI values of 3378 Twenty20 matches is of order 49803 $\times$ 1053. exhibits the set of all possible probabilities with which a $PI$ value (current state) succeeds a $3$-tuple of $PI$ values (previous state).
It is important to note that the system inevitably produces some negative $PI$ values, as it is designed to take into account the calculations until the end of the innings. In matches where the team batting second achieves its target, the final $PI$ value comes out to be negative. We will simply avoid such values for practical purposes since calculating the Pressure Index after a team has won the match does not make sense.
After the transition matrix has been obtained, any new sequence consisting of $s$ entries ($s\geq 3$) can be inserted to predict the outcome of the unknown $(s+1)$th entry of the sequence. The prediction process is not very simple. While a tuple of previous states is inserted, the model works in a way such that it searches for occurrences of the exact set of previous states (maintaining the order) in the transition matrix. If multiple current states are observed corresponding to a set of previous states with nonzero probabilities in the transition matrix, then the expected value of the current state is considered as the final outcome. However, if no occurrences of current states are found with nonzero probability for a set of previous states, then the sum of the states in the tuple is calculated and sets of previous states whose entries add up to the same sum are searched in the transition matrix. The expected value of the current state is then calculated following similar steps as described earlier. In either case, an outcome is obtained irrespective of whether the exact previous states are present in the observed dataset.
The following example shows the outcome when $(1.49,1.63,1.48)$ is entered as the new sequence. In cricketing terms, this translates to a situation where the Pressure Index of a team batting second have been recorded as $1.49$, $1.63$, and $1.48$ at the end of three consecutive overs during the second innings of a Twenty20 match. Given the target and the number of overs the innings is into, we are interested in predicting the probable Pressure Index of the team at the end of the next over.
The device provides a list of all possible Pressure Index values a team has exhibited in the training sample of $3378$ Twenty20 matches after having $PI$ values $1.49$, $1.63$, $1.48$ in that order. The probabilities corresponding to the occurrences of each possible $PI$ value are listed from the transition matrix, and then finally the expected outcome of the Pressure Index value is calculated from these probabilities.
Pressure Index Probability of occurrence Pressure Index Probability of occurrence
0.00 0.010310341 1.66 0.041241365
0.31 0.010310341 1.67 0.030931024
0.66 0.010310341 1.68 0.041241365
0.82 0.010310341 1.69 0.020620683
0.88 0.010310341 1.70 0.015465512
1.21 0.010310341 1.71 0.030931024
1.25 0.010310341 1.72 0.046396536
1.36 0.010310341 1.73 0.046396536
1.39 0.010310341 1.74 0.025775853
1.40 0.010310341 1.75 0.030931024
1.42 0.005155171 1.76 0.056706877
1.43 0.010310341 1.79 0.005155171
1.46 0.010310341 1.80 0.010310341
1.47 0.010310341 1.81 0.030931024
1.48 0.010310341 1.82 0.020620683
1.51 0.015465512 1.84 0.030931024
1.52 0.015465512 1.86 0.041241365
1.54 0.015465512 1.87 0.010310341
1.56 0.005155171 1.90 0.010310341
1.57 0.010310341 1.91 0.010310341
1.58 0.015465512 1.94 0.003402413
1.59 0.036086194 1.95 0.005155171
1.60 0.018867925 2.00 0.010310341
1.61 0.020620683 2.02 0.020620683
1.62 0.020620683 2.23 0.010310341
1.64 0.034333436 2.39 0.010310341
Expected Pressure Index outcome: 1.576342
Table 6.2.1 Expected outcomes of $PI$ values based on previous states
This means that if a team faces a situation where they have recorded Pressure Index values of $1.49$, $1.63$, $1.48$ in that order at the end of three consecutive overs during the second innings of a Twenty20 match, then the expected Pressure Index value at the end of the next over would be approximately $1.576342$.
6.3 Strategy for a controlled run chase using the Markov model
It has been observed from extensive studies that if the Pressure Curve during a run chase reaches values less than $0.5$, then the match is most likely to be won by the team batting second. From [Saikia et al., 2019], it is understood that
$\mathbb{P}(\text{match won by team batting second}\mid 0<PI<0.5)=1,$ (14)
and for the different subintervals of $PI$ values between $[0.5,3.5]$, viz. $[0.5,\lambda_{i_{1}}]$, $[\lambda_{i_{1}},\lambda_{i_{2}}]$, $\dots$, $[\lambda_{i_{m}},3.5]$, the values of
$\mathbb{P}(\text{match won by team batting second}\mid\lambda_{i}<PI<\lambda_{j})$
generally start with $1$ but keep on decreasing until they eventually reach 00. An immediate strategy for teams batting second in Twenty20 matches would be to try to bring down the Pressure Index as low as possible, and effectively below $0.5$, at some stage during the innings in order to guarantee a win.
With the transition matrix already obtained and the target known at the start of the innings, the expected $PI$ value at the end of the next over at any stage during the run chase can be calculated. This allows teams batting second to strategize accordingly and set mini targets for the next few overs to try to reduce their Pressure Index to below $0.5$.
For example, in the situation elaborated in Table 6.2.1, for the occurrence of a sequence of $PI$ values $1.49$, $1.63$, $1.48$, the expected Pressure Index at the end of the next over is $1.576342$ and the probability of this Pressure Index being less than $0.5$ is only about $2.0620682\%$. This is a situation where it is visibly difficult for the team batting second to immediately get a hold on the match.
(expected $\mathbb{P}(\text{expected }PI<0.5)=\mathbb{P}(PI=0)+\mathbb{P}(PI=0.31)=0.020620682$ (15)
6.3.1 Case Study: Using the model in a T20I match
Consider the Twenty20 match played between Pakistan and West Indies at National Stadium, Karachi on April 03, 2018. Pakistan were chasing a target of 154 and won the match by 8 wickets. The Pressure Index values of Pakistan’s batting innings are as follows.
Overs Cum runs Cum wickets Pressure Index Expected PI
1 5 0 1.038638 -
2 15 0 1.043422 -
3 32 0 0.990569 -
4 51 0 0.908084 1.090972
5 57 0 0.933522 0.823888
6 62 1 1.035470 0.816428
7 71 1 1.029399 0.919259
8 79 1 1.032575 0.941079
9 86 1 1.048083 1.052562
10 89 1 1.131503 1.13
11 98 1 1.113852 1.082961
12 107 1 1.084413 1.09
13 114 2 1.147530 1.1407
14 120 2 1.173178 1.177427
15 132 2 0.940734 0.947136
16 139 2 0.829457 0.826328
17 154 2 0 0.0734211
Table 6.3.1.1 Pressure Index of Pakistan at the end of every over
Note that the expected PI values have gradually converged towards the actual PI values, especially after the end of the 9th over. The strategy of determining the PI values after every over thus works well in this run chase.
6.3.2 A counter example for the strategy
Consider the Twenty20 match played between Bangladesh and Pakistan at Sher-E Bangla National Stadium, Mirpur on April 24, 2015. The hosts were chasing a target of 142 and won the match by 7 wickets. The Pressure Index values of Bangladesh’s batting innings are listed in the table below.
Overs Cum runs Cum wickets Pressure Index Expected PI
1 14 1 1.041062 -
2 16 2 1.081729 -
3 25 2 1.194013 -
4 31 2 1.224512 1.150613
5 34 3 1.294495 1.226884
6 44 3 1.371937 1.2118
7 50 3 1.414531 1.363112
8 54 3 1.495481 1.464337
9 65 3 1.458339 1.6028
10 75 3 1.429256 1.430688
11 89 3 1.286981 1.285
12 97 3 1.261276 1.265888
13 107 3 1.152069 1.153505
14 116 3 1.027661 1.03
15 123 3 0.929088 0.929653
16 138 3 0.252506 0.964372
17 143 3 0 0.2066667
Table 6.3.2.1 Pressure Index of Bangladesh at the end of every over
Bangladesh’s expected PI was gradually converging towards the actual PI values until it predicted $0.964372$ at the end of the 16th over, deviating significantly from the actual PI $0.252506$ which happened on the brink of a 15-run over. The probability of PI being less than $0.5$ at that stage was only $0.111152\%$. Incidentally, this was also the highest number of runs scored in an over among all 47 overs in that match. This brought down their $PI$ to much lesser than $0.5$ and they completed the run chase in the next over itself. This example shows how the indices can be monitored throughout the innings and how continuously revising the probabilities of keeping the Pressure Index within a suitable range helps a team remain in control during a run chase in Twenty20 cricket.
6.4 Building a strategy using phase-wise Markov models
Although 6.3 provides an insight into how Pressure Index values can be monitored throughout an innings to help build a strategy for a controlled run chase, it emerges as quite a general approach irrespective of a team’s specific patterns during run chases in certain situations. For a better understanding, one might be interested in studying the patterns of where a team’s Pressure Index lies at different stages during a successful run chase in Twenty20 cricket.
The $PI$ values at the end of the 6th, 10th and 16th overs of teams during several successful run chases (comprising 2773 out of the 3378 matches from our initial training set) are considered. The choice of these specific stages corresponds to the importance of the end of the powerplay (6th over), the halfway mark of the innings (10th over), and the start of the death overs (17th over). AIC, BIC and Kolmogorov-Smirnov tests have been performed over a few known distributions to determine which known distribution fits the data corresponding to $PI$ values the best.
Criteria Gamma Exponential Weibull
$AIC$ $-1513.574$ $7094.485$ $-1065.882$
$BIC$ $-1501.721$ $7100.412$ $-1054.029$
K-S test statistic $(D)$ $0.022463$ $0.50916$ $0.074801$
K-S p-value $0.1221$ $2.2\times 10^{-16}$ $6.905\times 10^{-14}$
Table 6.4.1 Results for $PI$ values at the end of the 6th over
Criteria Gamma Exponential Weibull
$AIC$ $1285.369$ $8229.105$ $1511.110$
$BIC$ $1297.222$ $8235.032$ $1522.962$
K-S test statistic $(D)$ $0.017787$ $0.47144$ $0.062958$
K-S p-value $0.345$ $2.2\times 10^{-16}$ $5.858\times 10^{-10}$
Table 6.4.2 Results for $PI$ values at the end of the 10th over
Criteria Gamma Exponential Weibull
$AIC$ $4931.218$ $9351.203$ $5006.789$
$BIC$ $4943.068$ $9357.128$ $5018.64$
K-S test statistic $(D)$ $0.025706$ $0.38125$ $0.036553$
K-S p-value $0.05169$ $2.2\times 10^{-16}$ $0.001233$
Table 6.4.3 Results for $PI$ values at the end of the 16th over
Lower AIC, BIC, and K-S test statistic values, along with higher p-values, denote that the gamma distribution fits well for the $PI$ values corresponding to each section of the data. This gives us an idea of how the $PI$ values at three different stages of an innings are distributed.
Figure 2: Density plots of phase-wise Pressure Index
Phases of the innings Estimated shape parameter Estimated rate parameter
Powerplay $68.72744$ $57.66806$
Middle overs $18.66357$ $11.17790$
Death overs $3.926073$ $1.985266$
Table 6.4.4 — Estimated parameters of Gamma distribution
Separate transition matrices corresponding to the three different stages of an innings can be constructed based on the process described in 6.2. Outcomes resulting from a tuple consisting of three $PI$ values will then be obtained based on the transition matrix corresponding to the specific section of the innings only, irrespective of $PI$ values occurring in the other two sections. Once the estimated $PI$ value of a team for a certain over of the innings is obtained, one can attempt to bring it down to $0.5$ as early as possible, according to the theory described in 6.3. The three-phase partition of the innings avoids the overgeneralization of teams’ strategies in run chases.
7 Direction of future research
Tracing Pressure curves ([Bhattacharjee and Lemmer, 2016]) and revising Pressure Index values with the help of the proposed Markov model helps to build a strategy to keep teams batting second in Twenty20 matches in control throughout the chase. However, several potential developments to the model can be considered as part of future research:
  1. i.
    Different teams have different patterns of a run chase. For example, teams with greater batting depths might score at a faster rate even after the fall of a wicket and still succeed, but the same strategy might not work for teams with lesser batting depths. Therefore, errors are expected to occur when using the transition probabilities corresponding to one team to predict Pressure Index values for a team with a different composition.
  2. ii.
    The proposed Markov model does not take into account any factors regarding venues, which might play a significant role in run chases in Twenty20 matches. Modifying the model to incorporate the effects of different venues might help to increase the accuracy when predicting Pressure Index values.
  3. iii.
    The model could be improved by considering the differences in the qualities and strengths of the teams involved in a match. Apart from venues, the ability of teams winning close contests, and the strength of opposition bowlers and teams, might affect a team’s performance. For example, a Pressure Index of say, $1.42$, at some stage of the innings against a stronger opposition may have a different impact compared to that against a weaker opposition. In matches between a top-ranked team and a lower-ranked team, the former is expected to capitalize towards the end of the match even from less favorable situations, primarily due to the significant difference in experience possessed by the two teams.
  4. iv.
    The formula (5.1.1) for Pressure Index restricts its usage to the second innings only, primarily because of the behavior of the Pressure curve in case of unsuccessful run chases. Certain variance stabilizing transformations (suggested by Saikia, Bhattacharjee and Mukherjee, $2019$), where $\frac{CRRR}{IRRR}$ is raised to a power $(\frac{CRRR}{IRRR})^{\alpha}$ for some suitable choice of $\alpha$, help to restrict the abrupt rise in $PI$ values at later stages of the innings during unsuccessful run chases. Such transformations can be applied in constructing a similar Markov model if one intends to use it for prediction purposes for teams batting first as well.
  5. v.
    The methodology used in this study could be extended to other formats of limited overs cricket, for example One-Day Internationals, with suitable modifications.